OpenStudy (anonymous):

Enter the elements of the following subsets of N: {x; x>7 ∧ x >13} My answear is {x; x>7 ∧ x >13} = (13, ∞) Is that answear correct? or how should I think?

OpenStudy (welshfella):

you are right values of x between 7 and 13 are not included because of the intersect symbol.

OpenStudy (zzr0ck3r):

But you wrote an interval, and it wants the elements in N. So we would list it like, \[\{13,14,15,16,...\}\] or \[\mathbb{N}-\{1,2,3,4,5,6,7,8,9,10,11,12,13\}\]

OpenStudy (zzr0ck3r):

interval notation implies continuum

OpenStudy (zzr0ck3r):

I think...

OpenStudy (zzr0ck3r):

or better \(\mathbb{N}\cap (13, \infty)\) ^^^thats the one

OpenStudy (welshfella):

yep I missed that

OpenStudy (zzr0ck3r):

I was thinking it would be ok if you used an interval as long as we said our universal set was the natural numbers, but I dont think that makes sense because the natural numbers do not form a system, not even a group. Maybe just make our own notation \[\mathbb{N}\cap (13, \infty)=(13,\infty)_{\mathbb{N}}\] :)